# AI-Face: A Million-Scale Demographically Annotated AI-Generated Face Dataset and Fairness Benchmark

### v1

@@ B.5 Details of Post-Processing
Random Brightness and Contrast: This post-processing method adjusts the brightness and contrast of the image within specified limits. By applying random brightness and contrast variations, it introduces changes in the illumination and contrast levels of the images. This evaluates detector’s robustness to different illumination conditions. The results in Fig. 6 are after we adjust brightness and contrast with shifting limits 0.2.
@@
Rotation: Rotates the image within a specified angle limit. This post-processing method is used to evaluate the detector’s robustness to changes in the orientation of objects within the image. The results in Fig. 6 are after we randomly rotate the image within a range of -45 to 45 degrees.
Figure B.1: Visualization of the image after different post-processing.
@@
Figure B.2: Robustness analysis in terms of utility and fairness under varying degrees of JPEG compression.
@@
Figure B.3: Robustness analysis in terms of utility and fairness under varying kernel sizes of Gaussian Blur.
@@
Figure B.4: Robustness analysis in terms of utility and fairness under varying degrees of Hue Saturation Value.
@@
Figure B.5: Robustness analysis in terms of utility and fairness under varying degrees of Rotations.
@@
Figure B.6: Robustness analysis in terms of utility and fairness under varying degrees of Brightness Contrast.
B.6 Additional Fairness Robustness Evaluation Results
@@
Fig. B.2 to Fig. B.6 demonstrate detectors’ robustness analysis in more detail as a function of different degrees of post-processing. Overall, ViT-B/16 [63] and UnivFD [67] show stronger robustness to various post-processing methods compared to other detection methods. Fairness-enhanced detectors do not have robustness against post-processing; this would be a direction for future studies to work on. Figure B.2 presents a detailed robustness analysis in terms of utility and fairness under varying degrees of JPEG compression. The utility of all detectors decreases as image quality is reduced. Among the detectors, UnivFD [67] exhibits the highest utility robustness, while ViT-B/16 [63] demonstrates the strongest fairness robustness. When considering Gaussian blur, ViT-B/16 stands out as the most robust detector in terms of utility, whereas EfficientB4 [62] shows the greatest robustness in terms of fairness. Against Hue Saturation Value adjustments, DAW-FDD [20] shows the strongest utility robustness, while UnivFD excels in fairness robustness. ViT-B/16 demonstrates superior robustness in both utility and fairness when facing rotations. For brightness contrast variations, DAG-FDD [20] is the most robust detector in terms of utility, while UnivFD once again shows superior robustness in terms of fairness.
Figure B.7: The utility-fairness trade-off of Fairness-enhanced methods.
B.7 Additional Fairness Generalization Evaluation Results
We conduct additional generalization experiments by using models trained on FF++ [2] to evaluate their generalization performance on our AI-Face test set. For these experiments, we utilize the trained weights and intra-domain performance metrics provided by [16]. Consequently, only the detectors with the pre-trained weights available from [16] are evaluated on our AI-Face test set. Results are shown in Table 6. We report the detailed performance on generation category subsets (i.e., Deepfake Videos, GANs, and DMs) and the overall performance on the whole test set. We observe that detectors exhibit significant performance degradation, approaching coin-toss performance when trained on FF++ and tested on our AI-Face test set. This suggests that detectors trained solely on one deepfake video dataset is not sufficient for detecting face images generated by current more advanced generation models. This also highlights the significance of our AI-Face dataset, which is extensive, diverse and comprehensive in generation methods to develop and evaluate existing AI face detectors. The lowest performance is observed with GANs, likely due to the higher variety of generation methods within this category. Conversely, performance on the Deepfake Videos subset is relatively better. This could be because, despite being different datasets, the deepfake videos may share similar generation methods, resulting in less variation in the artifacts present in the generated images.
Type Detector Intra- Domain (FF++) Cross-Domain (Ours w/o FF++) Test Subset Cross-Domain (Ours w/o FF++) Whole Test Set
Deepfake
Videos (3)
GANs (10) DMs (8)
AUC $F_{EO}$ AUC $F_{EO}$ AUC $F_{EO}$ AUC $F_{EO}$ AUC
Naive Xception [61] 96.370 104.961 77.766 139.963 58.228 110.977 78.622 101.194 72.649
EfficientB4 [62] 95.670 110.626 76.612 148.656 44.501 88.420 73.426 94.609 65.323
Frequency F3Net [64] 96.350 74.828 74.328 93.278 39.127 89.927 75.480 68.299 65.149
SPSL [65] 96.100 97.558 77.766 141.029 40.100 91.837 58.919 123.534 55.483
SRM [66] 95.760 60.855 74.900 89.903 57.572 73.209 77.954 57.775 72.474
Spatial UCF [16] 97.050 102.798 77.650 122.485 40.477 95.657 77.568 79.479 67.708
CORE [68] 96.380 69.717 76.506 95.727 45.549 79.161 82.112 72.424 70.662
Table 36: Fairness and utility cross-domain evaluation. All detectors are trained on FF++ (model weights and AUC on FF++ test set are from [25]) and evaluated on our Demographically Annotated AI-Face. The best-performing method is highlighted in red.
B.8 Full Results of Effect of Increasing the Size of Train Set
@@
In this section, we provide the full evaluation results tested under different sizes of train set, as shown from Table 37 to Table 40. Intersection $F_{EO}$ and AUC align with the results in Fig. 7 of the submitted manuscript.
Model Type
Naive Frequency Spatial Fairness-enhanced
Dataset Size Attribute Metric
Xception
 [61]
EfficientB4
 [62]
ViT-B/16
 [63]
F3Net
 [64]
SPSL
 [65]
SRM
 [66]
UCF
 [16]
UnivFD
 [67]
CORE
 [68]
DAW-FDD
 [20]
DAG-FDD
 [20]
PG-FDD
 [21]
20% Gender $F_{MEO}$ 1.725 0.366 0.863 1.523 0.916 1.818 0.652 0.369 2.196 1.657 1.549 1.428
$F_{DP}$ 1.944 2.239 2.618 2.083 2.305 1.586 2.811 2.317 1.543 1.823 2.269 2.106
$F_{OAE}$ 1.076 0.419 0.906 1.057 0.775 0.800 0.617 0.620 1.076 0.950 1.145 0.904
$F_{EO}$ 2.030 0.386 1.635 1.945 1.280 2.081 0.768 0.629 2.214 1.738 2.244 1.742
Race $F_{MEO}$ 14.155 11.039 10.108 13.887 12.235 15.231 11.756 14.625 16.804 16.116 12.021 12.645
$F_{DP}$ 23.488 20.018 22.360 23.285 22.782 22.998 22.994 22.628 25.752 23.457 23.093 22.572
$F_{OAE}$ 5.266 5.286 5.057 5.416 4.807 5.425 5.063 6.459 5.913 5.009 4.877 4.676
$F_{EO}$ 24.015 19.947 25.662 25.293 22.940 23.207 24.837 28.765 29.623 22.162 22.625 21.318
Age $F_{MEO}$ 6.766 5.613 5.335 7.254 5.765 6.506 8.761 5.411 7.208 5.948 5.672 5.769
$F_{DP}$ 5.086 5.581 6.666 5.089 5.561 4.659 6.170 6.073 4.556 5.080 5.291 5.337
$F_{OAE}$ 3.784 3.177 4.958 3.745 3.493 3.435 4.491 4.692 4.209 4.183 3.159 3.242
$F_{EO}$ 9.533 9.157 12.476 9.632 9.222 9.203 11.928 14.228 10.548 9.699 8.470 8.339
Intersection $F_{MEO}$ 17.912 12.056 14.781 17.613 14.966 19.221 14.360 17.533 20.977 19.466 15.288 15.734
$F_{DP}$ 25.299 22.237 23.053 25.005 23.895 25.807 23.863 23.563 27.720 25.542 24.273 24.374
$F_{OAE}$ 8.001 9.313 8.647 7.898 7.506 6.137 8.856 11.806 8.713 5.859 7.538 5.378
$F_{EO}$ 54.208 45.790 54.752 56.299 50.295 52.526 55.119 66.894 63.986 44.272 49.137 45.127
- ACC 95.175 94.292 93.972 94.913 95.084 95.534 95.249 90.810 94.835 94.996 95.243 95.602
AUC 98.620 99.055 98.765 98.284 98.851 98.026 98.728 96.404 98.237 98.403 98.731 98.533
AP 98.805 99.325 99.132 98.441 99.083 98.353 98.931 97.227 98.410 98.578 98.980 98.695
EER 5.563 5.208 6.267 6.142 5.292 5.489 4.933 10.001 6.169 5.696 5.424 5.148
Table 37: Detailed fairness and utility evaluation results on 20% training subset.
Model Type
Naive Frequency Spatial Fairness-enhanced
Dataset Size Attribute Metric
Xception
 [61]
EfficientB4
 [62]
ViT-B/16
 [63]
F3Net
 [64]
SPSL
 [65]
SRM
 [66]
UCF
 [16]
UnivFD
 [67]
CORE
 [68]
DAW-FDD
 [20]
DAG-FDD
 [20]
PG-FDD
 [21]
40% Gender $F_{MEO}$ 1.771 2.562 1.588 1.383 1.277 0.567 1.191 0.752 1.465 1.034 1.303 1.41
$F_{DP}$ 1.801 3.841 1.948 2.33 2.088 1.955 2.715 2.756 2.113 2.362 2.128 2.998
$F_{OAE}$ 0.908 0.439 0.881 1.117 0.661 0.078 1.281 0.625 0.971 0.811 0.793 1.193
$F_{EO}$ 1.799 3.236 1.809 2.034 1.415 1.095 2.286 1.023 1.796 1.474 1.51 2.263
Race $F_{MEO}$ 14.7 12.688 9.731 14.333 10.203 7.959 14.511 14.169 14.04 11.7 12.504 7.79
$F_{DP}$ 23.675 21.948 22.57 23.424 22.165 21.403 25.546 23.264 22.811 21.994 22.856 21.024
$F_{OAE}$ 5.079 6.318 4.774 4.986 4.49 3.571 5.708 6.282 5.222 3.819 4.448 4.043
$F_{EO}$ 23.443 17.727 22.852 22.917 20.787 17.734 30.682 30.21 22.288 17.342 20.633 18.703
Age $F_{MEO}$ 7.594 3.343 4.055 6.859 5.051 4.126 6.145 5.676 6.85 5.874 6.46 3.48
$F_{DP}$ 4.951 6.723 5.222 5.421 5.485 4.937 5.55 6.471 5.272 5.672 5.447 6.276
$F_{OAE}$ 3.873 1.951 2.928 3.709 3.057 2.589 3.747 4.713 3.447 3.655 3.689 3.158
$F_{EO}$ 9.596 9.58 8.258 8.736 8.461 8.457 9.256 14.995 8.374 9.236 9.119 8.222
Intersection $F_{MEO}$ 18.307 20.275 14.911 17.454 12.641 12.131 19.386 17.922 17.346 13.83 15.213 11.211
$F_{DP}$ 25.685 24.437 23.109 24.801 23.725 22.091 26.683 23.444 24.706 23.145 24.355 21.662
$F_{OAE}$ 5.814 10.624 8.96 5.964 5.707 6.691 10.527 11.615 5.902 4.63 5.04 6.579
$F_{EO}$ 48.63 41.478 49.564 47.936 44.562 37.57 63.173 66.301 49.171 34.392 43.13 40.847
- ACC 95.796 94.03 94.822 95.844 95.794 95.393 94.754 90.711 95.984 95.975 96.257 95.337
AUC 98.696 98.932 99.024 98.851 99.064 98.722 98.306 96.371 98.824 98.974 98.949 99.092
AP 98.778 99.269 99.31 98.959 99.236 98.968 98.588 97.224 98.984 99.139 99.035 99.318
EER 5.027 5.442 5.474 5.002 4.574 4.77 6.037 10.044 5.009 4.567 4.285 4.729
Table 38: Detailed fairness and utility evaluation results on 40% training subset.
Model Type
Naive Frequency Spatial Fairness-enhanced
Dataset Size Attribute Metric
Xception
 [61]
EfficientB4
 [62]
ViT-B/16
 [63]
F3Net
 [64]
SPSL
 [65]
SRM
 [66]
UCF
 [16]
UnivFD
 [67]
CORE
 [68]
DAW-FDD
 [20]
DAG-FDD
 [20]
PG-FDD
 [21]
60% Gender $F_{MEO}$ 0.801 0.489 2.239 1.576 1.787 1.179 0.745 0.539 0.907 1.512 1.527 0.408
$F_{DP}$ 2.899 2.747 4.223 1.996 1.638 3.062 2.596 2.509 2.291 3.01 1.88 2.857
$F_{OAE}$ 0.692 0.109 0.935 0.904 0.802 0.361 0.783 0.634 0.697 1.296 0.716 0.27
$F_{EO}$ 1.136 0.677 3.474 1.757 2.002 1.22 1.328 0.586 1.153 2.435 1.594 0.547
Race $F_{MEO}$ 8.652 16.885 6.328 13.433 16.19 6.243 9.96 14.482 14.243 9.849 14.223 5.96
$F_{DP}$ 21.794 26.205 14.609 23.519 21.498 18.874 20.947 23.469 24.031 20.453 23.547 15.86
$F_{OAE}$ 3.781 6.63 5.65 4.671 5.716 4.346 4.133 6.328 4.569 3.77 5.247 3.746
$F_{EO}$ 18.942 23.67 21.707 22.128 23.478 12.107 15.885 29.99 22.562 14.96 22.213 14.735
Age $F_{MEO}$ 5.047 5.153 5.719 6.155 4.154 3.81 4.71 5.243 5.789 5.512 3.699 3.553
$F_{DP}$ 6.012 4.411 7.664 5.157 5.456 6.02 6.042 6.245 5.444 7.926 5.023 6.488
$F_{OAE}$ 2.916 2.496 4.283 3.316 3.897 2.374 2.752 4.555 3.422 3.886 2.858 2.244
$F_{EO}$ 7.607 10.635 13.662 8.084 9.09 7.503 6.872 14.282 8.089 8.321 6.951 8.124
Intersection $F_{MEO}$ 10.466 25.134 7.982 16.425 17.532 10.693 12.44 17.613 16.374 12.417 17.272 9.574
$F_{DP}$ 22.891 27.88 18.106 25.118 24.338 20.236 22.678 23.819 25.063 22.277 25.459 18.176
$F_{OAE}$ 5.884 11.229 7.443 5.547 6.822 7.714 4.749 11.612 5.287 5.726 5.873 5.899
$F_{EO}$ 39.873 51.509 46.548 46.511 52.673 28.055 35.888 66.884 45.261 30.682 47.626 31.167
- ACC 96.505 93.931 93.612 96.221 95.676 96.51 96.567 90.882 96.009 95.025 96.332 96.488
AUC 98.97 98.828 98.536 99.075 99.102 99.236 99.026 96.461 99.189 99.003 99.354 99.401
AP 98.987 99.195 98.953 99.17 99.234 99.415 99.012 97.279 99.351 99.285 99.503 99.461
EER 3.829 6.004 6.668 4.322 4.314 3.248 3.592 9.875 4.351 5.072 3.882 3.583
Table 39: Detailed fairness and utility evaluation results on 60% training subset.
Model Type
Naive Frequency Spatial Fairness-enhanced
Dataset Size Attribute Metric
Xception
 [61]
EfficientB4
 [62]
ViT-B/16
 [63]
F3Net
 [64]
SPSL
 [65]
SRM
 [66]
UCF
 [16]
UnivFD
 [67]
CORE
 [68]
DAW-FDD
 [20]
DAG-FDD
 [20]
PG-FDD
 [21]
80% Gender $F_{MEO}$ 1.753 0.256 1.697 0.976 1.199 0.166 1.235 0.447 1.428 0.398 0.526 0.339
$F_{DP}$ 1.925 2.648 1.891 2.861 3.002 2.642 3.511 2.461 1.881 2.762 2.695 2.643
$F_{OAE}$ 0.943 0.002 0.988 0.280 0.474 0.316 0.737 0.596 0.665 0.214 0.408 0.172
$F_{EO}$ 1.893 0.364 1.910 1.214 1.489 0.218 1.237 0.467 1.495 0.522 0.788 0.384
Race $F_{MEO}$ 11.908 11.806 9.589 4.724 3.751 8.864 2.988 14.911 13.396 3.892 5.036 2.891
$F_{DP}$ 22.332 21.476 19.620 18.520 17.354 18.431 16.783 23.411 22.809 16.666 18.631 17.598
$F_{OAE}$ 4.298 6.127 4.724 3.890 3.573 4.030 2.667 6.322 4.970 3.282 4.350 2.731
$F_{EO}$ 18.793 19.118 20.637 10.997 9.458 14.889 10.966 29.610 22.226 9.621 13.148 8.090
Age $F_{MEO}$ 5.554 4.823 4.219 2.168 3.355 2.588 1.699 5.731 6.528 2.822 1.498 1.076
$F_{DP}$ 5.307 5.675 5.586 6.111 6.492 6.159 6.884 6.433 4.840 6.781 5.943 5.842
$F_{OAE}$ 3.001 2.397 3.365 1.221 1.905 1.389 0.832 4.916 3.150 2.718 1.133 0.744
$F_{EO}$ 7.274 8.476 8.710 4.026 8.252 5.533 5.139 15.746 8.380 7.114 4.174 2.835
Intersection $F_{MEO}$ 14.979 17.336 11.294 6.650 6.863 9.372 5.369 18.159 16.769 5.729 8.210 5.443
$F_{DP}$ 24.220 21.943 21.145 21.258 20.254 20.920 19.077 24.033 24.954 18.015 20.556 19.798
$F_{OAE}$ 5.025 10.608 7.908 6.697 5.709 6.343 5.118 11.541 5.558 5.760 7.955 4.583
$F_{EO}$ 40.744 44.028 45.684 27.249 22.360 32.012 24.504 66.750 46.492 21.906 29.401 17.687
- ACC 96.629 94.917 94.904 95.309 96.461 96.548 97.736 90.898 95.586 95.808 97.317 98.277
AUC 99.361 98.788 99.143 99.409 99.597 99.682 99.753 96.501 98.440 99.419 99.739 99.860
AP 99.429 99.051 99.403 99.523 99.653 99.765 99.801 97.308 98.562 99.589 99.817 99.874
EER 3.538 5.198 5.189 3.894 3.138 2.707 2.276 9.817 5.470 4.259 2.745 1.738
Table 40: Detailed fairness and utility evaluation results on 80% training subset.
B.9 Fairness and Utility Trade-off
Fig. B.7 presents the trade-offs between $F_{DP}$ on age and AUC of three fairness-enhanced methods. This is to analyze how well these methods balance optimizing utility and ensuring fairness in decision-making. 1) PG-FDD [21] achieves the best utility-fairness trade-off overall. It improves fairness without compromising the precision of utility, maintaining high accuracy in detection. For instance, PG-FDD achieves a higher AUC than DAW-FDD and DAG-FDD while maintaining comparable fairness metrics. 2) DAW-FDD [20] is sensitive to the hyperparameter that balances utility-fairness. For example, when its fairness approaches to zero, its utility also drops to a coin-tossing performance. This sensitivity can hinder practical deployment, as extensive tuning is required to optimize performance. 3) To ensure broader applicability and reliability, future fairness approaches should aim to minimize sensitivity to hyperparameter settings.
Appendix C Datasheet for AI-Face

### v3

@@ D.3 Details of Post-Processing
Random Brightness and Contrast: This post-processing method adjusts the brightness and contrast of the image within specified limits. By applying random brightness and contrast variations, it introduces changes in the illumination and contrast levels of the images. This evaluates detector’s robustness to different illumination conditions. The results in Fig. 6 are after we adjust brightness and contrast with shifting limits 0.4.
@@
Rotation: Rotates the image within a specified angle limit. This post-processing method is used to evaluate the detector’s robustness to changes in the orientation of objects within the image. The results in Fig. 6 are after we randomly rotate the image within a range of -30 to 30 degrees.
Figure D.2: Visualization of the image after different post-processing.
@@
Figure D.3: Robustness analysis in terms of utility and fairness under varying degrees of JPEG compression.
@@
Figure D.4: Robustness analysis in terms of utility and fairness under varying kernel sizes of Gaussian Blur.
@@
Figure D.5: Robustness analysis in terms of utility and fairness under varying degrees of Hue Saturation Value.
@@
Figure D.6: Robustness analysis in terms of utility and fairness under varying degrees of Rotations.
@@
Figure D.7: Robustness analysis in terms of utility and fairness under varying degrees of Brightness Contrast.
D.4 Additional Fairness Robustness Evaluation Results
@@
Fig. D.3 to Fig. D.7 demonstrate detectors’ robustness analysis in more detail as a function of different degrees of post-processing. Overall, ViT-B/16 [84] and UnivFD [88] show stronger robustness to various post-processing methods compared to other detection methods. Fairness-enhanced detectors do not have robustness against post-processing; this would be a direction for future studies to work on. Figure D.3 presents a detailed robustness analysis in terms of utility and fairness under varying degrees of JPEG compression. The utility of all detectors decreases as image quality is reduced. Among the detectors, ViT-B/16 [84] exhibits the highest utility robustness, ViT-B/16 [84] and UnivFD [88] both demonstrate the strongest fairness robustness. When considering Gaussian blur, ViT-B/16 again stands out as the most robust detector in terms of utility, whereas DAW-FDD [29] and UnivFD [88] show the great robustness in terms of fairness. Against Hue Saturation Value adjustments, SPSL [86] shows the strongest utility robustness, while the fairness of DAW-FDD [29] fluctuates less with different Hue Saturation Value adjustments. ViT-B/16 demonstrates superior robustness in both utility and fairness when facing rotations. For brightness contrast variations, SPSL [86] is the most robust detector in terms of utility, while UnivFD once again shows superior robustness in terms of fairness. Last, we can get the same conclusion from Fig. D.3 to Fig. D.7 as in the main manuscript, that post-processing clearly impairs detectors’ utility but does not necessarily make detectors more biased.
D.5 Full Results of Effect of Increasing the Size of Train Set
@@
In this section, we provide the full evaluation results tested under different sizes of train set, as shown from Table D.2 to Table D.5. Intersection $F_{EO}$ and AUC align with the results in Fig. 7 of the submitted manuscript.
D.6 Full Results of Effect of the Ratio of Real and Fake
In this section, we provide the full evaluation results tested under the train set with different ratios of real and fake, as shown from Table D.6 to Table D.8. Intersection $F_{EO}$ and AUC align with the results in Fig. 7 of the submitted manuscript.
D.7 Comparison Results with Foundation Model
In the Discussion (Section 5.3) of the main manuscript, we highlighted the potential of integrating foundation models (e.g., CLIP) into detector design as a strategy for mitigating bias. To explore this, we conducted a preliminary experiment by designing a detector using a frozen CLIP model combined with a trainable 3-layer MLP. This model was trained and tested on the AI-Face dataset. For comparison, we selected one representative detector from each model type: EfficientB4 [83], SPSL [86], UnivFD [88], and PG-FDD [30]. These four detectors’ results are consistent with those reported in Table 4. As shown in Table D.9, the CLIP+MLP detector demonstrates a clear advantage in both fairness and utility metrics, suggesting that foundation models hold significant promise for bias mitigation. For instance, its $F_{EO}$ score is 3.11% lower than the second-best method, PG-FDD, for the Skin Tone group, and 14.046% lower for the Intersection group.
Fairness Utiliy
Skin Tone Gender Age Intersection -
Method $F_{MEO}$ $F_{DP}$ $F_{OAE}$ $F_{EO}$ $F_{MEO}$ $F_{DP}$ $F_{OAE}$ $F_{EO}$ $F_{MEO}$ $F_{DP}$ $F_{OAE}$ $F_{EO}$ $F_{MEO}$ $F_{DP}$ $F_{OAE}$ $F_{EO}$ AUC ACC AP EER FPR
EfficientB4 5.385 1.725 1.487 5.863 8.300 6.184 4.377 11.062 6.796 11.849 2.856 10.300 17.586 8.607 8.461 25.114 98.611 94.203 99.542 6.689 20.066
SPSL 4.411 1.827 1.037 4.534 8.055 9.379 1.135 9.789 27.614 11.232 7.270 40.943 10.379 13.259 2.464 21.679 98.747 96.346 99.356 4.371 13.661
UnivFD 4.503 1.19 1.622 5.408 2.577 8.556 2.748 5.536 5.436 15.249 3.793 14.148 6.119 14.026 6.287 20.255 98.192 93.651 99.400 7.633 18.550
PG-FDD 3.190 1.252 1.071 3.702 6.465 9.746 0.882 9.115 14.804 10.467 5.009 29.585 9.578 14.697 3.062 18.348 99.172 96.174 99.694 4.961 10.971
CLIP+MLP 0.419 0.938 0.227 0.591 0.506 8.658 0.334 1.021 0.765 14.473 0.395 1.802 1.973 13.992 1.000 4.302 99.973 99.290 99.991 0.793 1.171
Table D.9: Fairness and utility performance of CLIP+MLP compared to representative detectors on the AI-Face dataset, highlighting the potential of foundation models for bias mitigation.
Appendix E Datasheet for AI-Face